A Bibliography of Publications about the Icon Programming Language

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

14 October 2017 Version 1.23

Abstract

This bibliography records books and other publications about the Icon programming language.

Title word cross-reference

2nd [HOP93].

’87 [BSK87, Wex87].

ACM [HOP93]. actors [Chr89]. Alternatives [FH92]. applications [Hin87]. approach [Jef99]. Assembly [Gri94]. Automatically [Slo95].

Based [PTTA99]. brief [Han93].

character [Gri80]. Chicon [WLL98]. Chinese [WLL98]. class [Wal89]. Co [WG83a]. Co-expressions [WG83a]. Collection [FH92]. command [FH83]. Compiler [Slo95, WG92]. computation [Gri94]. computer [BSK87]. Computing [Chr89]. Concurrent [Nil90a]. Conference [HOP93, BSK87]. Control [GN83].

Data [Gri89, Nil90b]. Defined [GN83]. Denotational [Gud92]. Design [GT93]. Directed [Gud92, Nil90a, Nil90b, OWG93, WG83b]. disparate [Hin87]. does [Gri94]. Driven [Chr89]. Dynamic [GT93].

Emacs [Mit84]. Evaluation [Gri82, Slo95, WG83b, Tra10]. Execution [JG94]. Experiences [Tra10]. exploratory [Jef99]. expression [Tra10]. Expressions [Gri82, WG83a]. extension [Hin87].

Fe [USE94]. Federal [BSK87]. First [Wal89]. First-class [Wal89]. Framework [JG94].

Garbage [FH92]. Generated [Slo95]. Generators [GHK81, Gri88, WG83b]. Germany [BSK87]. Glimpse [JM98]. Goal [Gud92, Nil90a, WG83b, Nil90b, OWG93]. Goal-Directed [Gud92, Nil90a, WG83b, Nil90b, OWG93]. Graphics [GJT97, GJT98, Jef94].

Handbook [Chr96]. Hashing [GT93]. held [BSK87]. High [Nil90a, USE94, FH83]. High-level [Nil90a, FH83]. History [HOP93, GG93]. HOPL [HOP93]. HOPL-II [HOP93]. Human [BSK87]. Human-computer [BSK87]. humanists [Cor90]. hybrid [Gri85].

Icon [Gri89, Shi96, Chr96, Cor90, FH92, GHK79, Gri80, GHK81, Gri82, GG83, Gri83, GG86, GG90, Gri90, GT93, GG93, Gri94, GJT97, GG97, GJT98, Han80, Han93, JG94, Jef94, JM98, LC86, Mit84, Nil90a, OG87, OWG93, PTTA99, Wal89, WG92, WG83a, WG83b]. Icon3 [Tra10]. Icon3-like [Tra10]. IFIP [BSK87]. II [HOP93]. Implementation [GG86, GT93, PTTA99, WG83b, OWG93]. Implementing [Gri83]. integrating [Hin87]. Integration [LC86]. INTERACT [BSK87]. Interaction [BSK87]. interpreter [OG87]. Interpreters [Wex87]. Interpretive [Wex87]. introduction [Han93].

Java [PTTA99]. Java-Based [PTTA99].

Jcon [PTTA99]. June [Wex87].

Language [Chr96, GG83, GG86, GO88, Gri89, GG90, Gri90, Gud92, Han80, JM98, WG92, WLL98, FH83, GHK79, GG93, GG97, Hin87, OG87, OWG93, Shi96, Gri94]. Languages [HOP93, USE94]. Level [USE94, FH83, Nil90a]. like [Tra10]. Logicon [LC86].

Management [Han80]. Manipulating [GO88]. Manipulation [WLL98]. mappings [Gri80]. Matching [Gri83, Nil90b]. Message [Chr89]. Mexico [USE94]. Minnesota [Wex87]. Monitoring [JG94, Jef99].

numerical [Gri94].

object [Hin87]. October [USE94].

operational [OWG93]. Operations [GN83]. Optimizing [WG92]. oriented [Hin87]. overview [GHK79].

Pattern [Gri83, Nil90b]. patterns [Wal89]. Paul [Wex87]. Icon [Gri85]. Portable [Han80]. Preprints [HOP93]. procedural [OWG93]. Proceedings [USE94, Wex87, BSK87]. Processing [Nil90a, Gri94]. Program [Jef99]. Programmer [GN83]. Programmer-Defined [GN83]. Programming [HOP93, Chr96, GG83, GG86, Gri88, GO88, Gri89, GG90, Gri90, Gri94, GJT97, Han80, Jef94, JMPP04, WG92, Cor90, FH83, GHK79, GG93, GG97, GJT98, OG87, Shi96]. Prolog [LC86].

Rebus [Gri85]. recursive [OG87].

relationship [Chr89]. Republic [BSK87]. Required [Gri94]. Rest [JM98]. Result [WG83c].

Santa [USE94]. Scanning [Gri90]. Second [BSK87]. Semantics [Gud92, OWG93]. September [BSK87]. Seque [GO88]. Sequences [GO88, WG83c, Nil90b]. Sets [GT93, Gri80]. SIGPLAN [HOP93, HOP93, Wex87]. SNOBOL4 [Gri85, Gri83]. SNOBOL4/Icon [Gri85]. Some [Gri94]. St. [Wex87]. Storage [Han80]. stream [Nil90b]. String [Gri90, Gri94]. structure [Gri94]. Structures [Gri89]. Stuttgart [BSK87]. subsystem [Mit84]. supports [Nil90b]. Symposium [USE94, Wex87]. System [Han80, Tra10].

Tables [GT93]. Techniques [Wex87]. Text [WLL98]. type [Nil90b].

unbounded [Nil90b]. Unicon [JMPP04]. University [BSK87]. UNIX [Mit84]. USA [USE94]. use [Gri80].

values [Nil90b]. Version [Jef94]. Very [USE94]. VHLL [USE94]. visualization [Jef99].

References

Bullinger:1987:HII

[BSK87] H.-J. Bullinger, B. Shackel, and K. Kornwachs, editors. Humancomputer Interaction — INTER-ACT ’87: proceedings of the Second IFIP Conference on Human-Computer Interaction, held at the University of Stuttgart, Federal Republic of Germany, 1–4 September, 1987. North-Holland, Amsterdam, The Netherlands, 1987. ISBN 0-444-70304-7. LCCN QA76.9.S88I325

Christopher:1989:MDC

[Chr89] Thomas W. Christopher. Message driven computing and its relationship to actors. ACM SIGPLAN Notices, 24(4):76–78, April 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org:
80/pubs/citations/proceedings/
plan/67386/p76-christopher/.

Christopher:1996:IPL

[Chr96] Thomas W. Christopher. Icon Programming Language Handbook. (none), (none), 1996. ISBN (none). 226 pp. LCCN (none).

URL http://toolsofcomputing.
com/IconHandbook/.

Corre:1990:IPH

[Cor90] Alan D. Corre. Icon programming for humanists. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1990. ISBN 0-13-450180-2. xiv + 157 pp. LCCN QA76.73.I19 C67 199.

Fraser:1983:HLP

[FH83] C. W. Fraser and D. R. Hanson. A high-level programming and command language. ACM SIGPLAN Notices, 18(6):212–219, June 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fernandez:1992:GCA

[FH92] Mary F. Fernandez and David R. Hanson. Garbage collection alternatives for Icon. Software—Practice and Experience, 22(8):659–672, August 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Griswold:1983:IPL

[GG83] Ralph E. Griswold and Madge T.

Griswold. The Icon Programming Language. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1983. ISBN 0-13-449777-5. xviii + 313 pp. LCCN QA76.73.I19 G74 1983.

Griswold:1986:IIP

[GG86] Ralph E. Griswold and Madge T.

Griswold. The Implementation of the Icon Programming Language. Princeton University Press, Princeton, NJ, USA, 1986. ISBN 0-691-08431-9. x + 336 pp. LCCN QA76.73.I19 G76 1986.

US$39.50. URL http://www.cs.
arizona.edu/icon/ibsale.htm.

Griswold:1990:IPL

[GG90] Ralph E. Griswold and Madge T.

Griswold. The Icon Programming Language. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, second edition, 1990. ISBN 0-13-447889-4. xv + 367 pp. LCCN QA76.73.I19 G74 1990.

Griswold:1993:HIP

[GG93] Ralph E. Griswold and Madge T.

Griswold. History of the Icon programming language. ACM SIGPLAN Notices, 28(3):53–68, March 1993. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org:
80/pubs/citations/proceedings/
plan/154766/p53-griswold/.

Griswold:1997:IPL

[GG97] Ralph E. Griswold and Madge T. Griswold. The Icon programming language. Peer-to-Peer Communications, San Jose, CA, USA, third edition, 1997. ISBN 1-57398-001-3. xx + 386 pp. LCCN QA76.73.I19G74 1997. US$34.95. URL http://www.

                                        cs.arizona.edu/icon/lb3.htm.

Griswold:1979:IPL

[GHK79] R. E. Griswold, D. R. Hanson, and J. T. Korb. The Icon programming language: an overview. ACM SIG-PLAN Notices, 14(4):18–31, April 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Griswold:1981:GI

[GHK81] Ralph E. Griswold, David R. Hanson, and John T. Korb. Generators in Icon. ACM Transactions on Programming Languages and Systems, 3 (2):144–161, April 1981. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Griswold:1997:GPI

[GJT97] Ralph E. Griswold, Clinton L. Jeffery, and Gregg M. Townsend. Graphics Programming in Icon. Peer-to-Peer Communications, San Jose, CA, USA, 1997. ISBN 1-57398-009-9. ???? pp.

Griswold:1998:GPI

[GJT98] Ralph E. Griswold, Clinton L. Jeffery, and Gregg M. Townsend. Graphics programming in Icon. Peerto-Peer Communications, San Jose, CA, USA, 1998. ISBN T385.G765 1998. xiv + 504 pp. LCCN 1-57398-009-9.

Griswold:1983:PCO

[GN83] R. E. Griswold and Michael Novak.

Programmer-defined control operations. The Computer Journal, 26(2): 175–183, May 1983. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Griswold:1988:SPL

[GO88] Ralph E. Griswold and Janalee O’Bagy. Seque: a programming language for manipulating sequences. Computer Languages, 13(1):13–22, ???? 1988. CODEN COLADA. ISSN 0096-0551.

Griswold:1980:UCS

[Gri80] R. E. Griswold. The use of character sets and character mappings in Icon. The Computer Journal, 23(2): 107–114, May 1980. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Griswold:1982:EEI

[Gri82] Ralph E. Griswold. The evaluation of expressions in Icon. ACM Transactions on Programming Languages and Systems, 4(4):563–584, October 1982. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Griswold:1983:ISP

[Gri83] Ralph E. Griswold. Implementing Snobol4 pattern matching in Icon. Computer Languages, 8(2):77– 92, ???? 1983. CODEN COLADA. ISSN 0096-0551.

Griswold:1985:RSI

[Gri85] R. E. Griswold. Rebus — a SNOBOL4/Icon hybrid. ACM SIG-PLAN Notices, 20(2):7–16, February 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Griswold:1988:PG

[Gri88] R. E. Griswold. Programming with generators. The Computer Journal, 31(3):220–228, June 1988. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Griswold:1989:DSI

[Gri89] Ralph E. Griswold. Data structures in the icon programming language. Computing Systems, 2(4): 339–365, Fall 1989. CODEN CM-SYE2. ISSN 0895-6340.

Griswold:1990:SSI

[Gri90] R. E. Griswold. String scanning in the Icon programming language. The Computer Journal, 33(2):98– 107, April 1990. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Griswold:1994:SAR

[Gri94] Ralph E. Griswold. Some assembly required: The Icon Programming Language: Icon does string and structure processing and numerical computation. Byte Magazine, 19(5): 193–??, May 1994. CODEN BYT-EDJ. ISSN 0360-5280.

Griswold:1993:DID

[GT93] William G. Griswold and Gregg M. Townsend. The design and implementation of dynamic hashing for sets and tables in Icon. Software— Practice and Experience, 23(4):351– 367, April 1993. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gudeman:1992:DSG

[Gud92] David A. Gudeman. Denotational semantics of a goal-directed language. ACM Transactions on Programming Languages and Systems, 14(1):107–125, January 1992. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
104659.html.

Hanson:1980:PSM

[Han80] David R. Hanson. A portable storage management system for the Icon programming language. Software— Practice and Experience, 10(6):489– 500, June 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hanson:1993:BII

[Han93] David R. Hanson. A brief introduction to Icon. ACM SIGPLAN Notices, 28(3):359–360, March 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://

www.acm.org:80/pubs/citations/
proceedings/plan/154766/p359-
hanson/.

Hinckley:1987:OOE

[Hin87] K. Hinckley. An object oriented extension language for integrating disparate applications. In Bullinger et al. [BSK87], pages 529– 533. ISBN 0-444-70304-7. LCCN QA76.9.S88I325 1987.

ACM:1993:ASH

[HOP93] ACM SIGPLAN HOPL-II. 2nd ACM SIGPLAN History of Programming Languages Conference (Preprints), volume 28(3). ACM Press, New York, NY, USA, March 1993. CODEN SINODQ. ISBN ???? ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN ????

Jeffery:1994:GPI

[Jef94] Clinton L. Jeffery. Graphics programming in Icon version 9. In USENIX Association [USE94], pages 157–168. ISBN 1-880446-65-0. LCCN QA76.7 .U74 1994.

Jeffery:1999:PMV

[Jef99] Clinton L. Jeffery. Program monitoring and visualization: an exploratory approach. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 0-387-98644-8. xiv + 209 pp. LCCN QA76.65.J45 1999.

Jeffery:1994:FEM

[JG94] Clinton L. Jeffery and Ralph E. Griswold. A framework for execution monitoring in Icon. Software —Practice and Experience, 24(11): 1025–1049, November 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jeffery:1998:GIL

[JM98] Clinton Jeffery and Shamim Mohamed. A glimpse of Icon: A language for the rest of us. Linux Journal, 51:??, July 1998. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Jeffery:2004:PU

[JMPP04] Clinton L. Jeffery, Shamim Mohamed, Ray Pereda, and Robert Parlett. Programming with Unicon. ????, ????, 2004. ISBN ???? xiii + 371 pp. LCCN

????  URL http://unicon.
sourceforge.net/book/ub.pdf.

Lapalme:1986:LIP

[LC86] Guy Lapalme and Suzanne Chapleau. Logicon: An integration of Prolog into Icon. Software—Practice and Experience, 16(10):925–944, October 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mitchell:1984:ISU

[Mit84] William H. Mitchell. An Icon subsystem for UNIX emacs. Technical report TR 84-8, Dept. of Computer Science, The University of Arizona, Tucson, AZ, USA, 1984. 17 pp.

Nilsen:1990:HLG

[Nil90a] Kelvin Nilsen. High-level goaldirected concurrent processing in Icon. Software—Practice and Experience, 20(12):1273–1290, December 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Nilsen:1990:SDT

[Nil90b] Kelvin Nilsen. A stream data type that supports goal-directed pattern matching on unbounded sequences of values. Computer Languages, 15(1): 41–54, ???? 1990. CODEN CO-LADA. ISSN 0096-0551.

OBagy:1987:RII

[OG87] J. O’Bagy and R. E. Griswold. A recursive interpreter for the Icon programming language. ACM SIGPLAN Notices, 22(7):138–149, July 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.org:
80/pubs/citations/proceedings/
plan/29650/p138-o_bagy/.

OBagy:1993:OSI

[OWG93] Janalee O’Bagy, Kenneth Walker, and Ralph E. Griswold. An operational semantics for Icon: implementation of a procedural goal-directed language. Computer Languages, 18 (4):217–239, ???? 1993. CODEN COLADA. ISSN 0096-0551.

Proebsting:1999:JJB

[PTTA99] Todd Proebsting, Gregg Townsend, Denise Todd, and Bob Alexander. Jcon: A Java-based Icon implementation. Version 2 of Jcon is an essentially complete implementation of Icon, including graphics and large integers. A few minor features are missing — mostly things like chdir() that cannot be done in

Java., 1999. URL http://www.cs.
arizona.edu/icon/jcon/.

Shipman:1996:IPL

[Shi96] John Shipman. The icon programming language. This tutorial on Icon is available as a World-Wide Web document., April 11,

1996. URL http://www.nmt.edu/
tcc/help/lang/icon/.

Sloane:1995:EAG

[Slo95] Anthony M. Sloane. An evaluation of an automatically generated compiler. ACM Transactions on Programming Languages and Systems, 17(5):691–703, September 1995. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
213980.html.

Tratt:2010:EIL

[Tra10] Laurence Tratt. Experiences with an Icon3-like expression evaluation system. ACM SIGPLAN Notices, 45 (12):73–80, December 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

USENIX:1994:PUS

[USE94] USENIX Association, editor. Proceedings of the USENIX Symposium on Very High Level Languages (VHLL): October 26–28, 1994, Santa Fe, New Mexico, USA. USENIX, Berkeley, CA, USA, 1994. ISBN 1-880446-65-0. LCCN QA76.7 .U74 1994.

Walker:1989:FPI

[Wal89] Kenneth W. Walker. First-class patterns for Icon. Computer Languages, 14(3):153–163, ???? 1989. CODEN COLADA. ISSN 0096-0551.

Wexelblat:1987:IIT

[Wex87] Richard L. Wexelblat, editor. Proceedings of the SIGPLAN ’87 Symposium on Interpreters and Interpretive Techniques, St. Paul, Minnesota, June 24–26, 1987. ACM Press, New York, NY, USA, 1987. ISBN 0-89791-235-7. LCCN QA76.7 .S54 v.22:7. US$23.00. SIGPLAN Notices, v. 22, no. 7 (July 1987).

Wampler:1983:CI

[WG83a] S. B. Wampler and R. E. Griswold. Co-expressions in Icon. The Computer Journal, 26(1):72–78, February 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Wampler:1983:IGG

[WG83b] Stephen B. Wampler and Ralph E. Griswold. The implementation of generators and goal-directed evaluation in Icon. Software—Practice and Experience, 13(6):495–518, June 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wampler:1983:RS

[WG83c] Stephen B. Wampler and Ralph E. Griswold. Result sequences. Computer Languages, 8(1):1–14, ???? 1983. CODEN COLADA. ISSN 0096-0551.

Walker:1992:OCI

[WG92] Kenneth Walker and Ralph E. Griswold. An optimizing compiler for the Icon programming language. Software—Practice and Experience, 22 (8):637–657, August 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wong:1998:CCT

[WLL98] Kam-Fai Wong, Vincent Y. Lum, and Wai-Ip Lam. Chicon — A Chinese text manipulation language. Software—Practice and Experience, 28(7):681–701, July 1998.

CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

URL http://www3.interscience.
wiley.com/cgi-bin/abstract?ID=
1804;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?ID=
1804&PLACEBO=IE.pdf.